Using Kirchhoff's junction rule, determine which

Part A

Using Kirchhoff's junction rule, determine which one of the following four diagrams would be the appropriate one to use for solving this problem.

Using Kirchhoff's junction rule, determine which



Part B
Calculate the current I2 flowing in emf source E2.
Answer #1

According to Kirchhoffs junction rule, the sum of currents towards the junction is equal to the sum of currents away from thApply Kirchhoffs loop rule for the top loop in the figure D, Apply Kirchhoffs loop rule for the bottom loop in the figure D
Answer #2

Solution

Part a)

For Figure A, applying Kirchhoff's junction rule we have that:

I_{2}+I_{2}=I_{1}-I_{2}\, \, \, \, \, \, \, \, \Rightarrow \, \, \, \, \, \, \, \, \, \, I_{2}=I_{1}

Therefore, figure A is not correct.

For figure B, applying Kirchhoff's junction rule we have that:

I_{1}+I_{2}=I_{2}-I_{1}+I_{2}\, \, \, \, \, \, \, \, \Rightarrow \, \, \, \, \, \, \, \, \, \, I_{1}=I_{2}-I_{1}

therefore, figure B is not correct

For figure C, applying Kirchhoff's junction rule we have that:

I_{2}=I_{2}+I_{2}-I_{1}\, \, \, \, \, \, \, \, \Rightarrow \, \, \, \, \, \, \, \, \, \, I_{1}=I_{2}

therefore, figure C is not correct.

For figure D, applying Kirchhoff's junction rule we have that:

I_{1}-I_{2}+I_{2}=I_{1}\, \, \, \, \, \, \, \, \Rightarrow \, \, \, \, \, \, \, \, \, \, I_{1}=I_{1}

Therefore, the correct figure is figure D:

2 2

Part b)

と 2 2 2

Applying Kirchhoff's loop rule at loop 1 we have:

-V_{1}+I_{1}R+V_{2}=0\, \, \, \, \, \, \, \, \, \, \, \, \, \, \, \, \, \, \, (1)

but:

V_{1}=\varepsilon _{1}-I_{1}r_{1}\, \, \, \, \, \, \, \, \, \, \, \, \, \, \, \, \, \, \, (2)

and

と2

Replacing (2) and (3) into equation (1):

-\varepsilon _{1}+I_{1}r_{1}+I_{1}R+\varepsilon _{2}-I_{2}r_{2}=0\, \, \, \, \, \, \, \, \, \, \, \, \, \, \, \, \, \, \, (4)

or

I_{1}\left ( R+r_{1} \right )-I_{2}r_{2}=\varepsilon _{1}-\varepsilon _{2}\, \, \, \, \, \, \, \, \, \, \, \, \, \, \, \, \, \, \, (5)

Simillarly, applying Kirchhoff's loop rule at loop 2 we have:

-V_{2}+(I_{1}+I_{2})R_{b}=0\, \, \, \, \, \, \, \, \, \, \, \, \, \, \, \, \, \, \, (6)

Replacing (3) into equation (6):

I_{1}R_{b}+I_{2}(R_{b}+r_{2})=\varepsilon _{2}\, \, \, \, \, \, \, \, \, \, \, \, \, \, \, \, \, \, \, (7)

Grouping equations (5) and (7) we have:

\left\{\begin{matrix} I_{1}\left ( R+r_{1} \right )-I_{2}r_{2}=\varepsilon _{1}-\varepsilon _{2} & & & \\ I_{1}R_{b}+I_{2}(R_{b}+r_{2})=\varepsilon _{2}& & & \end{matrix}\right.

Solving the system for I_{2} we have:

{\color{Blue} I_{2}=\frac{\varepsilon _{2}(1+R_{b})-\varepsilon _{1}R_{b}}{r_{2}R_{B}+\left ( R_{b}+r_{2} \right )\left ( R+r_{1} \right )}}
Answer #3

Part-A The junction rule states that the current coming towards the junction is always equal to the current going away from tPart-B Use the Kirchhoffs loop law in upper loop of the circuit in figure-A -E2-1R-Ii-I,r 0 Rearrange the above-mentioned eq